OpenStudy (anonymous):

Why is it when you integrate 1/(1+u^2) du, you would get tan^-1+C? How did it come to that?

OpenStudy (anonymous):

let u=tantheta du=...??

OpenStudy (anonymous):

it actually came from the fact that d/dx arctan(x) = 1/(x^2+1)

OpenStudy (anonymous):

@sourwing Is that just a known rule or something?

OpenStudy (anonymous):

I don't know what you meant by known rule, but derivative of arctan(x) is derived using implicit differentiation

OpenStudy (zarkon):

and the inverse function theorem

OpenStudy (anonymous):

If possible, could someone the whole process of getting the derivative of arctanx?

OpenStudy (anonymous):

if y = arctan(x), then x = tan(y), then using implicit differentiation, 1 = sec^2(y) dy/dx dy/dx = 1/cos^2(x) = 1/(x^2 + 1) Because d/dx (arctanx) = 1/(x^2 + 1), this is why the integral of 1/(x^2+1) is arctan(x).

OpenStudy (zarkon):

\[\sec^2(y)=\tan^2(y)+1=x^2+1\]

OpenStudy (zarkon):

also, you need the inverse function theorem in order to make sure \[\frac{dy}{dx}\] exists

OpenStudy (anonymous):

yeah, I left a lot of details out. But the goal is i'm trying to explain that why integral of arctan(x) is 1/(x^2+1)

OpenStudy (zarkon):

you mean the derivative of arctan(x) is 1/(x^2+1)

OpenStudy (anonymous):

that too :DD

OpenStudy (kainui):

\[\sin^2 \theta + \cos ^2 \theta =1 \] divide both sides by cos^2(theta) \[\tan^2 \theta + 1 = \sec^2 \theta\] In the integral, we notice if we pick u=tan(theta) we will get a nice simplification. Take the derivative of u with respect to theta to get: \[\frac{ du }{ d \theta } = \sec^2 \theta\] and multiply both sides by d theta to plug into your integral: \[\int\limits_{}^{}\frac{ du }{ u^2+1 } = \int\limits_{}^{}\frac{ \sec^2 \theta d \theta }{ \tan^2\theta+1 }=\int\limits_{}^{} d \theta\]

OpenStudy (anonymous):

Thank you everyone for your help.

OpenStudy (kainui):

Of course from here since we know u=tan(theta) then theta must be arctan(u). This is the better method since you can easily integrate other similar things. Suppose you have to integrate: \[\int\limits_{}^{} \frac{ dx }{ x^2 + 9 }\] then what do you do? Do you take the derivative of stuff by hacking in the dark? No. You choose: x=3tan(theta) since when it plugs in you can easily see you'll get: \[9\tan^2 \theta + 9 = 9 \sec^2 \theta\]